OpenStudy (mw647):

Find the center and radius of (x + 8)2 + (y + 4)2 = 49. (8, 4); 7 (–8, –4); 7 (–4, 8); 7

OpenStudy (paxpolaris):

\(\left( x-h \right)^2 +\left( y-k \right)^2=r^2\) has a center at \(\left( h,k \right)\) and radius \(r\)

OpenStudy (paxpolaris):

(–8, –4); 7

OpenStudy (mw647):

Thank you!